OpenStudy (anonymous):

Come on, someone please help me! I have asked the same question three times and no one is helping. I just need someone to check my answer! please help!!

OpenStudy (anonymous):

wheres your question?

OpenStudy (anonymous):

no question here ....

OpenStudy (anonymous):

its right here: Assume a 115-pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 : Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2: Write an equation, in slope intercept form, to represent this data. Part 3 : How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned. Answer: Part 1: Place a point at 12 x and 36 y (12,36). Then place another point at 18 x and 54 y (18, 54). Then, draw a line through both points. The slope is 3 ( 12 - 18 / 36 - 54) this is the rate of how many calories are burns a minute. Part 2: y = 3x Part 3: If we already know what the x represents, then all you have to do is plug in the numbers to find the solution. For example, If x = 35 then y = 3(35) or 105 calories burned.

OpenStudy (anonymous):

does that sound good, please let me know ASAP!!

OpenStudy (anonymous):

Part 1)  slope = (54 - 36) / (18-12) = 3